OpenStudy (anonymous):

The average of the two fractions are? 9 2/3 and 11 3/4

OpenStudy (vishal_kothari):

10.7083

OpenStudy (anonymous):

10.7083 as a fraction...?

OpenStudy (anonymous):

\[\large\frac{\frac{29}{3}+\frac{47}{4}}{2}\] \[\frac{\frac{116}{12}+\frac{141}{12}}{2}=\frac{\frac{116+141}{12}}{2}=\frac{257}{24}\]

OpenStudy (anonymous):

Dockworker - That equals to... 10 17/24?

OpenStudy (anonymous):

you can do simple division to convert to improper fraction

OpenStudy (anonymous):